OpenStudy (anonymous):

What is the domain and range of the graph of 2/(x^2-2x-3)

OpenStudy (shamim):

i m trying to solve it

OpenStudy (shamim):

\[y=\frac{ 2 }{ x ^{2}-2x-3 }\]

OpenStudy (shamim):

\[or, y=\frac{ 2 }{ x ^{2}-3x+x-3 }\]

OpenStudy (shamim):

\[or, y=\frac{ 2 }{ x(x-3)+(x-3) }\]

OpenStudy (shamim):

\[or, y=\frac{ 2 }{ (x+1)(x-3) }\]

OpenStudy (shamim):

actually the value of x is called domain n the value of y is called range

OpenStudy (shamim):

n the denominator never b 0

OpenStudy (shamim):

because the value of y is became undefined by the value of 0 denominator

OpenStudy (shamim):

so the domain will b any real number accept -1 and 3

OpenStudy (shamim):

because\[x+1\neq0\]

OpenStudy (shamim):

\[or, x \neq-1\]

OpenStudy (shamim):

again\[x-3\neq0\]

OpenStudy (shamim):

\[or, x \neq3\]

OpenStudy (shamim):

and y can get any real number

OpenStudy (shamim):

so range will b any real number

OpenStudy (shamim):

if u need more explanation then ask me plz

OpenStudy (anonymous):

Thanks!

OpenStudy (shamim):

u r most welcome